# Optimal Control for Steady Circulation of a Diffusion Process via Spectral Decomposition of Fokker–Planck Equation

### 18 Mar 2026 v1

@@ III-A Control Objective
The control objective includes fast convergence to the stationary distribution $\rho_{\mathrm{s}}$ of the uncontrolled system. Here, we consider that the system to have converged to the stationary state when $\rho$ is sufficiently close to $\rho_{\mathrm{s}}$ within a finite time. The stationary flux $\tilde{\bm{J}}_{\mathrm{s}}$ in the stationary state with $\rho=\rho_{\mathrm{s}}$ should satisfy the divergence-free condition for all $\bm{x}\in\Omega$, i.e.,
@@
In general, $\nabla\cdot(\rho_{\mathrm{s}}\nabla\alpha)$ does not vanish but it is sufficient that the condition (16) is satisfied when $u_{1}=0$, because $u_{1}$ becomes practically zero after the sufficient convergence of $\rho$ to $\rho_{\mathrm{s}}$. Meanwhile, the condition (17) is always satisfied because the following relationship holds for any smooth function $\phi$:
$\nabla\cdot\left(\nabla^{\bot}\phi\right)=0.$ (18)
Therefore, when $u_{1}=0$, $\rho_{\mathrm{s}}$ is still a stationary distribution of the FPE (10) even for $u_{2}\neq 0$.
@@ III-B Spectral Dimensionality Reduction
We consider the eigenfunction expansion of the PDF $\rho$ using the eigenfunctions $\{v_{m}\}$ of the linear operator $\mathscr{L}^{*}$, i.e.,
@@
$\partial_{t}\rho(\bm{x},t)=\mathscr{L}^{*}\rho(\bm{x},t)+u_{1}(t)\zeta_{1}(\bm{x},t)+u_{2}(t)\zeta_{2}(\bm{x},t).$ (23)
The dynamics of each coefficient $c_{m}$ can then be described by the following ordinary differential equation (ODE):
@@ III-C Optimization
To achieve the control objective, we introduce a stage cost functional $\tilde{\mathcal{L}}_{1}$ for the PDF, which is described using a weight $Q_{1}>0$ as follows:
$\tilde{\mathcal{L}}_{1}[\rho]={}$ $\frac{Q_{1}}{2}\left|\rho(\cdot,t)-\rho_{\mathrm{s}}\right|_{L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)}^{2},$ (34)
where $\left|\cdot\right|_{L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)}$ represents the $L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)$ norm. In addition, we introduce a stage cost functional $\tilde{\mathcal{L}}_{2}$ for the flux rotation, which is described using a weight $Q_{2}>0$ as follows:
$\tilde{\mathcal{L}}_{2}[\omega]={}$ $\frac{Q_{2}}{2}\left|\omega(\cdot,t)-\omega_{\mathrm{d}}\right|_{L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)}^{2}.$ (35)
We also introduce a terminal cost functional $\tilde{\varphi}$ for the PDF and flux rotation using a similar formulation as
$\tilde{\varphi}[\rho(\cdot,t_{\mathrm{f}}),\omega(\cdot,t_{\mathrm{f}})]={}$ $\frac{Q_{\mathrm{f}}}{2}\left|\rho(\cdot,t_{\mathrm{f}})-\rho_{\mathrm{s}}\right|_{L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)}^{2}$ (36)
@@
Hence, the objective functional $\tilde{\mathcal{J}}$ is formulated as
$\tilde{\mathcal{J}}[u_{1},u_{2}]={}$ $\tilde{\varphi}[\rho(\cdot,t_{\mathrm{f}}),\omega(\cdot,t_{\mathrm{f}})]$ (39)
$+\int_{t_{0}}^{t_{\mathrm{f}}}\tilde{\mathcal{L}}_{1}[\rho]+\tilde{\mathcal{L}}_{2}[\omega]+\mathcal{L}_{3}(u_{1},u_{2})dt.$~~ (40)
The proposed optimal control problem satisfies the standard conditions ensuring the existence of an optimal control, namely coercivity and weak lower semicontinuity of the cost functional together with continuity of the control-to-state mapping [14, 12].
Next, we reformulate the objective functional by using the eigenfunction expansion (20) of the PDF $\rho$. The stage cost for the PDF can be approximated by a function $\mathcal{L}_{1}$ of the coefficient vector $\bm{c}$ as
$\tilde{\mathcal{L}}_{1}[\rho]$ (41)
$\approx{}$ $\mathcal{L}_{1}(\bm{c})$ (42)
@@
where the $m$-th component of the vector $\bm{c}_{\mathrm{s}}\in\mathbb{R}^{M}$ is obtained as $c_{\mathrm{s}}^{(m)}=\left\langle\rho_{\mathrm{s}},v_{m}\right\rangle_{L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)}$ and $\left\|\cdot\right\|$ represents the Euclid norm of a $M$-dimensional vector. The stage cost for the flux rotation can also be approximated by a function $\mathcal{L}_{2}$ of $\bm{c}$ and $u_{1,2}$ by calculating the flux $\tilde{\bm{J}}$ using the eigenfunction expansion (20), which can be written as
$\tilde{\mathcal{L}}_{2}[\omega]$ (45)
$\approx{}$ $\mathcal{L}_{2}(\bm{c},u_{1},u_{2})$ (46)
$={}$ $\frac{Q_{2}}{2}\left\langle-\nabla^{\bot}\cdot\tilde{\bm{J}}(\cdot,t)-\omega_{\mathrm{d}},v_{k}\right\rangle_{L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)}$ (47)
$={}$ $\frac{Q_{2}}{2}\left\|(\bm{A}_{1}+u_{1}(t)\bm{A}_{2}+u_{2}(t)\bm{A}_{3})\bm{c}-\bm{d}\right\|^{2}.$ (48)
@@
respectively, because the flux $\tilde{\bm{J}}$ can be represented as
$\tilde{\bm{J}}$={} $-\sum_{m=0}^{M-1}c_{m}v_{m}\left(\nabla V+u_{1}\nabla\alpha+u_{2}\frac{1}{\rho_{\mathrm{s}}}\nabla^{\bot}\phi\right)$ (53)
@@
$\tilde{\varphi}[\rho(\cdot,t_{\mathrm{f}}),\omega(\cdot,t_{\mathrm{f}})]$ (55)
$\approx{}$ $\varphi(\bm{c}(t_{\mathrm{f}}),u_{2}(t_{\mathrm{f}}))$ (56)
@@
$\tilde{\mathcal{J}}[u_{1},u_{2}]$ (58)
$\approx{}$ $\mathcal{J}[u_{1},u_{2}]$ (59)
$={}$ $\varphi(\bm{c}(t_{\mathrm{f}}),u_{2}(t_{\mathrm{f}}))$ (60)
$+\int_{t_{0}}^{t_{\mathrm{f}}}\mathcal{L}_{1}(\bm{c})+\mathcal{L}_{2}(\bm{c},u_{1},u_{2})+\mathcal{L}_{3}(u_{1},u_{2})dt.$ (61)
The objective functional includes the product of $\bm{c}$ and $u_{1,2}$ but is still positive definite.
The optimal control problem can be formulated with a Lagrange multiplier vector $\bm{\mu}(t)\in\mathbb{R}^{M}$ as
$\min_{u_{1},u_{2}}\quad\mathcal{J}[u_{1},u_{2}]+\int_{t_{0}}^{t_{\mathrm{f}}}\bm{\mu}(t)^{\top}\left(\bm{f}(\bm{c},u_{1},u_{2})-\dot{\bm{c}}(t)\right)dt$ (62)
@@
given that the initial condition at time $t_{0}$ for the state $\bm{c}$ satisfies
@@
where $\bm{c}_{0}$ is the initial state corresponding to the initial PDF $\rho_{0}$ and its $m$-th component is obtained as $(c_{0})^{(m)}=\left\langle\rho_{0},v_{m}\right\rangle_{L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)}$.
@@
$H(\bm{c},u_{1},u_{2},\bm{\mu})={}$ $\mathcal{L}_{1}(\bm{c})+\mathcal{L}_{2}(\bm{c},u_{1},u_{2})+\mathcal{L}_{3}(u_{1},u_{2})$ (65)
@@
Since it is difficult to obtain the optimal solution analytically, we numerically obtained it using the MATLAB optimization solver fminunc. At each step of the optimization, the state equation (32) is first numerically integrated from $t=t_{0}$ to $t=t_{\mathrm{f}}$ using the Euler method with the initial condition (64), and then the ODE for $\bm{\mu}$ given by (67) is solved from $t=t_{\mathrm{f}}$ to $t=t_{0}$ with the terminal condition (68), to evaluate the value of the objective functional $\mathcal{J}$. Finally, we obtained a local optimal solution that satisfies the stationarity condition of the Hamiltonian with respect to the inputs, i.e., $\partial H/\partial u_{1,2}=0$ for all $t\in[t_{0},t_{\mathrm{f}}]$. Although this optimization problem is not convex, the formulated optimization problem requires low computational cost owing to the dimensionality reduction of the original FPE.
@@ IV RESULTS
Figure 5: The behavior of the flux obtained from the particles under control at the final time on the xyxy plane. The length of the arrows represents the flux speed.

### 25 Mar 2026 v2

@@ III-A Control Objective
The control objective includes fast convergence to the stationary distribution $\rho_{\mathrm{s}}$ of the uncontrolled system. Here, we consider that the system to have converged to the stationary state when $\rho$ is sufficiently close to $\rho_{\mathrm{s}}$ within a finite time. The stationary flux $\tilde{\bm{J}}_{\mathrm{s}}$ in the stationary state with $\rho=\rho_{\mathrm{s}}$ should satisfy the following condition for all $\bm{x}\in\Omega$, i.e.,
@@
In general, $\nabla\cdot(\rho_{\mathrm{s}}\nabla\alpha)$ does not vanish but it is sufficient that the condition (16) is satisfied when $u_{1}=0$, because $u_{1}$ becomes practically zero after the sufficient convergence of $\rho$ to $\rho_{\mathrm{s}}$. Meanwhile, the condition (17) is always satisfied because $\nabla\cdot\left(\nabla^{\bot}\phi\right)=0$ holds for any smooth function $\phi$: Therefore, when $u_{1}=0$, $\rho_{\mathrm{s}}$ is still a stationary distribution of the FPE (10) even for $u_{2}\neq 0$.
@@ III-B Spectral Dimensionality Reduction
We consider the eigenfunction expansion of the PDF $\rho$ using the eigenfunctions $\{v_{m}\}$ of the linear operator $\mathscr{L}^{*}$ as
@@
$\partial_{t}\rho={}$ $\mathscr{L}^{*}\rho+u_{1}\zeta_{1}+u_{2}\zeta_{2}$ (22)
$\coloneqq{}$ $\mathcal{F}\{\rho,u_{1},u_{2}\},$ (23)
where $\mathcal{F}$ describes the dynamics of the controlled FPE. The dynamics of each coefficient $c_{m}$ can then be described by the following ordinary differential equation (ODE):
@@ III-C Optimization
To achieve the control objective, we introduce a stage cost functional $\mathcal{L}_{1}$ for the PDF, which is described using a weight $Q_{1}>0$ as follows:
$\mathcal{L}_{1}[\rho]={}$ $\frac{Q_{1}}{2}\left|\rho(\cdot,t)-\rho_{\mathrm{s}}\right|_{L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)}^{2},$ (34)
where $\left|\cdot\right|_{L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)}$ represents the $L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)$ norm. In addition, we introduce a stage cost functional $\mathcal{L}_{2}$ for the flux rotation, which is described using a weight $Q_{2}>0$ as follows:
$\mathcal{L}_{2}[\omega]={}$ $\frac{Q_{2}}{2}\left|\omega(\cdot,t)-\omega_{\mathrm{d}}\right|_{L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)}^{2}$ (35)
$={}$ $\frac{Q_{2}}{2}\left|\nabla\times\tilde{\bm{J}}(\cdot,t)-\omega_{\mathrm{d}}\right|_{L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)}^{2}.$ (36)
We also introduce a terminal cost functional $\varphi$ for the PDF and flux rotation using a similar formulation as
$\varphi[\rho(\cdot,t_{\mathrm{f}}),\omega(\cdot,t_{\mathrm{f}})]={}$ $\frac{Q_{\mathrm{f}}}{2}\left|\rho(\cdot,t_{\mathrm{f}})-\rho_{\mathrm{s}}\right|_{L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)}^{2}$ (37)
@@
Hence, the objective functional $\mathcal{J}$ is represented as
$\mathcal{J}[u_{1},u_{2}]={}$ $\varphi[\rho(\cdot,t_{\mathrm{f}}),\omega(\cdot,t_{\mathrm{f}})]$ (40)
$+\int_{t_{0}}^{t_{\mathrm{f}}}\mathcal{L}_{1}[\rho]+\mathcal{L}_{2}[\omega]+\mathcal{L}_{3}(u_{1},u_{2})dt.$~~ (41)
The optimal control problem is formulated as
$\min_{u_{1},u_{2}}$ $\quad\mathcal{J}[u_{1},u_{2}]$ (42)
$\mathrm{s.t.}$ $\quad\partial_{t}\rho=\mathcal{F}\{\rho,u_{1},u_{2}\},$ (43)
$\quad\rho(t_{0})=\rho_{0},$ (44)
where $\rho_{0}$ is the initial PDF at $t_{0}$. The proposed optimal control problem satisfies the standard conditions ensuring the existence of an optimal control, namely coercivity and weak lower semicontinuity of the cost functional together with continuity of the control-to-state mapping [14, 12].
Next, we reformulate the objective functional by using the eigenfunction expansion (18) of the PDF $\rho$. The stage cost for the PDF can be approximated by a function $\tilde{\mathcal{L}}_{1}$ of the coefficient vector $\bm{c}$ as
$\mathcal{L}_{1}[\rho]$ (45)
$\approx{}$ $\tilde{\mathcal{L}}_{1}(\bm{c})$ (46)
@@
where the $m$-th component of the vector $\bm{c}_{\mathrm{s}}\in\mathbb{R}^{M}$ is obtained as $c_{\mathrm{s}}^{(m)}=\left\langle\rho_{\mathrm{s}},v_{m}\right\rangle_{L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)}$ and $\left\|\cdot\right\|$ represents the Euclid norm of a $M$-dimensional vector. The stage cost for the flux rotation can also be approximated by a function $\tilde{\mathcal{L}}_{2}$ of $\bm{c}$ and $u_{1,2}$ by calculating the flux $\tilde{\bm{J}}$ using the eigenfunction expansion (18), which can be written as
$\mathcal{L}_{2}[\omega]\approx{}$ $\tilde{\mathcal{L}}_{2}(\bm{c},u_{1},u_{2})$ (49)
$={}$ $\frac{Q_{2}}{2}\left\|(\bm{A}_{1}+u_{1}(t)\bm{A}_{2}+u_{2}(t)\bm{A}_{3})\bm{c}-\bm{d}\right\|^{2}.$~ (50)
@@
respectively, because the flux $\tilde{\bm{J}}$ can be approximated as
$\tilde{\bm{J}}\approx{}$ $-\sum_{m=0}^{M-1}c_{m}v_{m}\left(\nabla V+u_{1}\nabla\alpha+u_{2}\frac{1}{\rho_{\mathrm{s}}}\nabla^{\bot}\phi\right)$ (55)
@@
$\varphi[\rho(\cdot,t_{\mathrm{f}}),\omega(\cdot,t_{\mathrm{f}})]$ (57)
$\approx{}$ $\tilde{\varphi}(\bm{c}(t_{\mathrm{f}}),u_{2}(t_{\mathrm{f}}))$ (58)
@@
$\mathcal{J}[u_{1},u_{2}]$ (60)
$\approx{}$ $\tilde{\mathcal{J}}[u_{1},u_{2}]$ (61)
$={}$ $\tilde{\varphi}(\bm{c}(t_{\mathrm{f}}),u_{2}(t_{\mathrm{f}}))$ (62)
$+\int_{t_{0}}^{t_{\mathrm{f}}}\tilde{\mathcal{L}}_{1}(\bm{c})+\tilde{\mathcal{L}}_{2}(\bm{c},u_{1},u_{2})+\mathcal{L}_{3}(u_{1},u_{2})dt.$ (63)
The objective functional includes the product of $\bm{c}$ and $u_{1,2}$ but is still coercive.
Thus, the low-dimensional optimal control problem can be formulated with a Lagrange multiplier vector $\bm{\mu}(t)\in\mathbb{R}^{M}$ as
$\min_{u_{1},u_{2}}\quad\tilde{\mathcal{J}}[u_{1},u_{2}]+\int_{t_{0}}^{t_{\mathrm{f}}}\bm{\mu}(t)^{\top}\left(\bm{f}(\bm{c},u_{1},u_{2})-\dot{\bm{c}}(t)\right)dt$ (64)
@@
given that the initial condition at $t_{0}$ for the state $\bm{c}$ satisfies
@@
where $\bm{c}_{0}$ is the initial state corresponding to $\rho_{0}$ and its $m$-th component is obtained as $(c_{0})^{(m)}=\left\langle\rho_{0},v_{m}\right\rangle_{L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)}$.
Theorem 1
In the $M\to\infty$ limit, any optimal solution of the approximated problem (64) converges to an optimal solution of the original problem (42).
Proof:
Since the eigenfunction expansion of $\rho$ up to m=M1m=M-1 converges to $\rho$ as $M\to\infty$ in $L^{2}\left(\Omega;\rho_{\mathrm{s}}^{-1}\right)$ space [18], the cost function $\tilde{\mathcal{L}}_{1}$ converges to the cost functional $\mathcal{L}_{1}$. Since the flux $\tilde{\bm{J}}$ can be expanded as Eq. (55) and $\omega(\bm{x},t)=-\nabla^{\bot}\cdot\tilde{\bm{J}}(\bm{x},t)$, the cost function $\tilde{\mathcal{L}}_{2}$ converges to the cost functional $\mathcal{L}_{2}$. Similarly, the cost function $\tilde{\varphi}$ converges to the cost functional $\varphi$. Therefore, the value of $\tilde{\mathcal{J}}$ converges to that of $\mathcal{J}$. Since $\tilde{\mathcal{J}}$ is coercive independently of $M$, any optimal solution of (64) remains bounded for all $M$ and converges to an optimal solution of (42). ∎
@@
$H(\bm{c},u_{1},u_{2},\bm{\mu})={}$ $\tilde{\mathcal{L}}_{1}(\bm{c})+\tilde{\mathcal{L}}_{2}(\bm{c},u_{1},u_{2})+\mathcal{L}_{3}(u_{1},u_{2})$ (67)
@@
Since it is difficult to obtain the optimal solution analytically, we numerically obtained it using the MATLAB optimization solver fminunc. At each step of the optimization, the state equation (32) is first numerically integrated from $t=t_{0}$ to $t=t_{\mathrm{f}}$ using the Euler method with the initial condition (66), and then the ODE for $\bm{\mu}$ given by (69) is solved from $t=t_{\mathrm{f}}$ to $t=t_{0}$ with the terminal condition (70), to evaluate the value of the objective functional $\tilde{\mathcal{J}}$. Finally, we obtained a local optimal solution that satisfies the stationarity condition of the Hamiltonian with respect to the inputs, i.e., $\partial H/\partial u_{1,2}=0$ for all $t\in[t_{0},t_{\mathrm{f}}]$. Although this optimization problem is not convex, the formulated optimization problem requires low computational cost owing to the dimensionality reduction of the original problem.
@@ IV RESULTS
Figure 5: The behavior of the flux obtained from the particles under the optimal control at the final time on the xyxy plane. The length of the arrows represents the flux speed.